# Contextual Document Embeddings

### v2

@@ 1 Introduction
For contextual training, we aim to build a notion of neighboring documents directly into the contrastive learning process. We propose a method that uses on fast query-document clustering to produce a group of neighbors for each training batch. Each update for training is constructed purely from neighboring documents to ensure that embeddings can distinguish documents even in the most challenging contexts.
@@ 2 Related Work
Much research has focused on the effect of hard negatives on the performance of contrastive learning methods Chen et al. (2020); Qu et al. (2021); Robinson et al. (2021); Wang et al. (2023). (Zhang & Stratos, 2021) observe that harder negatives provide a better approximation of the overall cross-entropy loss, but do not consider batch-level optimizations for negative selection. Hofstätter et al. (2021) cluster queries before training and show that this improves performance. Sachidananda et al. (2023) also consider contrastive batch sampling as a global optimization problem, but do not apply their technique to state-of-the-art transformer-based text embedding models. (Ma et al., 2024) use a clustering algorithm to partition a dataset into several sub-datasets, but train a different model on each sub-dataset. Our training algorithm aims to find the hardest possible batches to train text embedding model.
@@ 4.1 Contextual Training with Adversarial Contrastive Learning
To avoid a situation where each batch contains a large number of false negatives, we compute an equivalence class: $S(q,d)=\{d^{\prime}\in\mathcal{D}\mid f(q,d^{\prime})\geq f(q,d)\}$ for some surrogate scoring function $f$. At training time, we alter the partition function for $d$ so that it no longer includes the elements of $S(q,d)$, which are not definitively negative examples:
@@ 4.2 Contextual Document Embedding (CDE)
Contextualization can also be added directly to the archiecture. Taking inspiration from sparse vector retrieval which uses corpus statistics to determine the form of the embedding, we modify the encoders to have access to the corpus itself, i.e. $\phi(d;{\cal D})$ and $\psi(d;{\cal D})$. This effectively augments the biencoder model to give it the ability to contextualize documents directly.
@@ 5 Experimental Setup
For the large setting, we use the best settings found via small experiments. We train a single model on sequences of length $512$ with $512$ contextual documents, evaluating on the full MTEB benchmark (Muennighoff et al., 2022). This includes tasks from retrieval as well as tasks like classification, clustering, and reranking.

### v3

@@ 1 Introduction
For contextual training, we aim to build a notion of neighboring documents directly into the contrastive learning process. We propose a method that uses fast query-document clustering to produce a group of neighbors for each training batch. Each update for training is constructed purely from neighboring documents to ensure that embeddings can distinguish documents even in the most challenging contexts.
@@ 2 Related Work
Much research has focused on the effect of hard negatives on the performance of contrastive learning methods Chen et al. (2020); Qu et al. (2021); Robinson et al. (2021); Wang et al. (2023). (Zhang & Stratos, 2021) observe that harder negatives provide a better approximation of the overall cross-entropy loss, but do not consider batch-level optimizations for negative selection. Hofstätter et al. (2021) cluster queries before training and show that this improves performance. Sachidananda et al. (2023) also consider contrastive batch sampling as a global optimization problem, but do not apply their technique to state-of-the-art transformer-based text embedding models. (Ma et al., 2024) use a clustering algorithm to partition a dataset into several sub-datasets, but train a different model on each sub-dataset. Solatorio (2024) also use a pre-trained model to address the problem of in-batch false negatives from randomly sampled batches. Our training algorithm aims to find the hardest possible high-quality batches to train text embedding models.
@@ 4.1 Contextual Training with Adversarial Contrastive Learning
To avoid a situation where each batch contains a large number of false negatives, we compute an equivalence class: $S(q,d)=\{d^{\prime}\in\mathcal{D}\mid f(q,d^{\prime})\geq f(q,d)+{\epsilon}\}$ for some surrogate scoring function $f$ and boundary term ${\epsilon}$. At training time, we alter the partition function for $d$ so that it no longer includes the elements of $S(q,d)$, which are not definitively negative examples:
@@ 4.2 Contextual Document Embedding (CDE)
Contextualization can also be added directly to the architnhlltivvffrrhgcfvbjgnurgkhbclnlvecture. Taking inspiration from sparse vector retrieval which uses corpus statistics to determine the form of the embedding, we modify the encoders to have access to the corpus itself, i.e. $\phi(d;{\cal D})$ and $\psi(d;{\cal D})$. This effectively augments the biencoder model to give it the ability to contextualize documents directly.
@@ 5 Experimental Setup
For the large setting, we use the best settings found via small experiments. We train a single model on sequences of length $512$ with $512$ contextual documents, evaluating on the full MTEB benchmark (Muennighoff et al., 2022). This includes tasks from retrieval as well as tasks like classification, clustering, and reranking. We compare our model’s performance to the top small-size (under 250M parameters) models on MTEB (Nussbaum et al., 2024; Xiao et al., 2024; Solatorio, 2024; Li et al., 2023).